Radiation of a certain wavelength causes electrons with a maximum kinetic energy of 0.95 eV to be ejected from a metal whose work function is 2.55 eV. What will be the maximum kinetic energy (in eV) with which this same radiation ejects electrons from another metal whose work function is 1.81 eV?First metal Second metal

In the Compton effect, an X-ray photon of wavelength 0.16 nm is incident on a stationary electron. Upon collision with the electron, the scattered X-ray photon continues to travel in the same direction as the incident photon. What is the wavelength λ' of the scattered photon?
Answer #1

Given
maximum kinetic energy =0.95 eV
work function is 2.55 eV
The energy of the incident Photon is
hf=KEmax + Wo
=0.95+2.55
=3.5 ev
In the second case by rearranging
KEmax=hf-Wo
Here work function is 1.81 eV
=3.5-1.81
=1.69ev